OpenStudy (anonymous):

Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function. 5. Create your own equation written in standard form.

OpenStudy (anonymous):

give me a medal ill give you a dollar (: (:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes i already told you this is a function as each number increases by 1 and then 2 and then 3 and so on

OpenStudy (kewlgeek555):

@croslali I am also having a problem with this question. Well, not really. I said that it was a function well - this is what I put: I think this sequence can be a function. I am not sure how to explain this function, though. Hm. F(n) F(1) = 5 F(2) = f(1) + 1 F(3) = f(2) + 2 F(4) = f(3) + 3 F(5) = f(4) + 4 And so on and so on? Should I take out the whole recursive formula thingie?

OpenStudy (anonymous):

sorry im dead sleepy .. w ill think of it later now im going bye .. sorry man .. i wish i had the energy to do this

OpenStudy (anonymous):

sleep?

OpenStudy (anonymous):

were r u

OpenStudy (anonymous):

in london?

OpenStudy (anonymous):

its day here lol

OpenStudy (anonymous):

QATAR

OpenStudy (anonymous):

woah!cool

OpenStudy (anonymous):

u r awesome then

OpenStudy (anonymous):

on hols

OpenStudy (anonymous):

bye then

OpenStudy (kewlgeek555):

Well thanks anyways @croslali